OpenStudy (anonymous):

if a number is added to the numerator of 3/4 and twice as much is added to the denominator, the result is 4/7. find the number

OpenStudy (anonymous):

5 is the number

OpenStudy (anonymous):

3/4 + x/2x =4/7

OpenStudy (anonymous):

x/2x= 4/7 - 3/4

OpenStudy (anonymous):

x/2x= - 5/28

OpenStudy (anonymous):

2x(-5)=x(28)

OpenStudy (anonymous):

-10x= 28x

OpenStudy (anonymous):

unless x is 0, I did something wrong

OpenStudy (gw2011):

Let n=the number to be added (3+n)/(4+2n)=4/7 3+n=(4/7)(4+2n) 7(3+n)=4(4+2n) 21+7n=16+8n 21-16=8n-7n 5=n To verify this answer you get: (3+5)/(4+10)=8/14=4/7

OpenStudy (anonymous):

Oh... thanks!